STT 5) Cos/130°) 6) ta
Answer #1

los(138) = (0) ☺ ) los ( 13 18 cas l ñ-50 5.) -6157 - - les 155 Cos (138) -0.64 278 tan (57) = tan (7-7) -- tani = كم Han 157